# Rescaled Influence Functions: Accurate Data Attribution in High Dimension

### v1

@@ Abstract
How does the training data affect a model’s behavior? This is the question we seek to answer with data attribution. The leading practical approaches to data attribution are based on influence functions (IF). IFs utilize a first-order Taylor approximation to efficiently predict the effect of removing a set of samples from the training set without retraining the model, and are used in a wide variety of machine learning applications. However, especially in the high-dimensional regime (# params $\geq\Omega($# samples))))), they are often imprecise and tend to underestimate the effect of sample removals, even for simple models such as logistic regression. We present rescaled influence functions (RIF), a new tool for data attribution which can be used as a drop-in replacement for influence functions, with little computational overhead but significant improvement in accuracy. We compare IF and RIF on a range of real-world datasets, showing that RIFs offer significantly better predictions in practice, and present a theoretical analysis explaining this improvement. Finally, we present a simple class of data poisoning attacks that would fool IF-based detections but would be detected by RIF.
@@ 1 Introduction
The ability to quickly and accurately predict a leave-$T$-out effect, or to search for subsets producing a large leave-out effect, unlocks extensive capabilities from classical statistical inference to modern machine learning. For example, the jackknife, leave-$k$-out cross-validation, and bootstrap are all widely used to quantify uncertainty and estimate generalization error or confidence intervals, and all rely on the ability to quickly estimate leave-$T$-out effects Efr (92); GSL+ (19); Jae (72). Machine learning has seen an explosion of applications of data attribution, for dataset curation KL (17); KATL (19), explainability KATL (19); GBA+ (23), crafting and detection of data poisoning attacks EIC+ (25); KSL (22); SS (19), machine unlearning SAKS (21); GGHVDM (19); IASCZ (21), credit attribution JDW+ (19); GZ (19), bias detection BAHAZ (19), and more.
Ascertaining the ground truth leave-$T$-out effect in general requires a full retrain of a model for each $T$ of interest, which is computationally intractable in all but the simplest settings. Consequently, approximations to the leave-$T$-out effect are widely used. Key desiderata for such approximations are (1) accuracy, (2) computational efficiency even for large-scale models, and (3) additivity: the predicted effect of removing $T$ should be the sum of predicted effects of removing each element of $T$ individually. Additivity enables another important capability: search for the subset $T$ of a given size with the greatest predicted effect according to a given metric, by taking the $k$ training data points with largest predicted leave-one-out effects BGM (20); IPE+ (22).
@@
We introduce a simple and fast-to-compute modification of the influence function, which we term the rescaled influence function (RIF). RIFs improve accuracy by incorporating a limited amount of higher-order information about the change in model parameters from sample removal, but retain the additivity and in many settings also the computational efficiency of IFs. We show via experiments and theoretical analysis that RIFs are accurate for data attribution in overparameterized models where IFs struggle. Like IFs, RIFs are model and task agnostic, meaning that they can be applied to any empirical risk minimization-based training method with smooth losses, and they can estimate the leave-$T$-out effect according to any (smooth) measure of change to model parameters. We therefore advocate using RIFs as a drop-in replacement for IFs across data attribution applications.
@@ 1.1 Influence Functions, Newton Steps, and Rescaled Influence Functions
We now introduce our main contribution, the rescaled influence function, formally. Suppose that $\{(x_{i},y_{i})\}_{i\in[n]}$ is a training data set, $\Theta\subseteq\mathbb{R}^{d}$ is a class of models, and $\ell(x,y,\theta)$ is a twice-differentiable loss function; $\ell$ may include a regularizer. For simplicity, we imagine that $\ell$ is convex, although the definition of RIFs can be extended to the non-convex case. Let $\hat{\theta}=\arg\min_{\theta\in\Theta}\sum_{i\leq n}\ell(x_{i},y_{i},\theta)$ be the empirical loss minimizer (or, in the non-convex setting, any local minimum of the empirical loss).
@@
$\text{IF}_{i}=-\left[\frac{d}{dw_{i}}\cdot\hat{\theta}(w)\right]\Big{|}_{w=\mathbf{1}}=H^{-1}\cdot\nabla\ell(x_{i},y_{i},\hat{\theta})\,.$
@@
$\hat{\theta}_{\text{NS},T}=\hat{\theta}-H_{[n]\setminus T}^{-1}\left(\sum_{i\notin T}\nabla\ell(x_{i},y_{i},\hat{\theta})\right)=\hat{\theta}+H_{[n]\setminus T}^{-1}\left(\sum_{i\in T}\nabla\ell(x_{i},y_{i},\hat{\theta})\right)\,.$
@@
The accuracy of the NS approximation comes at significant cost, since each fresh $T$ requires a Hessian inversion, and additivity is lost. The RIF recovers additivity and much of the computational efficiency of IF, but retains much of the accuracy of the NS approximation. For sample $i\in[n]$, let $\textup{RIF}_{i}$ be the NS approximation to the leave-$i$-out effect, given by $\textup{RIF}_{i}=H_{[n]\setminus\{i\}}^{-1}\cdot\nabla\ell_{i}(x_{i},y_{i},\hat{\theta})$. Then for $T\subseteq[n]$, we define the RIF approximation to the leave-$T$-out effect to be
@@
The computational overhead of RIF compared to IF depends in general on the cost of computing the $n$ leave-one-out Hessian inversions – once these are obtained, no fresh Hessian inversion is needed to compute $\hat{\theta}_{\textup{RIF},T}$ for any $T$. RIF is especially attractive in generalized linear models, where $\textup{RIF}_{i}$ can be obtained from $\text{IF}_{i}$ by multiplying by a rescale factor $(1-h_{i})^{-1}$, where $h_{i}$ is a (generalized) leverage score associated to the $i$-th sample, which can be computed via a single matrix-vector product with $H^{-1}$. Thus, for generalized linear models, no additional Hessian inversion is needed. For example, in logistic regression, the formula for $\textup{RIF}_{i}$ uses the rescaling $(1-h_{i})^{-1}$, where $h_{i}=\hat{y_{i}}(1-\hat{y_{i}})\cdot x_{i}^{\top}H^{-1}x_{i}$; here $\hat{y_{i}}\in[0,1]$ is the logistic predicted label of the $i$-th sample according to $\hat{\theta}$.
Beyond generalized linear models, whenever each sample makes a low-rank contribution the Hessian, the $n$ leave-one-out Hessian inversions can be computed quickly via the Sherman-Morrison/Woodbury formula. In all of our experiments, the running time overhead to compute RIF compared to IF is negligible.
@@ 3 Theoretical Results
We think of each $\ell_{i}$ as a per-sample loss from the $i$-th sample in an underlying training set, although we do not actually need to assume such a training set underlies the optimization problem. Let ${\mathbf{g}}_{i}:=\nabla\ell_{i}(\hat{{\bm{\theta}}})$ and ${\mathbf{H}}_{i}:=\nabla^{2}\ell_{i}(\hat{{\bm{\theta}}})$ denote the gradient and Hessian of the $i$th sample at the solution $\hat{{\bm{\theta}}}$, and define the total Hessian ${\mathbf{H}}:=\sum_{i=1}^{n}{\mathbf{H}}_{i}$.
@@
Ultimately, we use IF/RIF/NS to estimate the change to $f(\hat{{\bm{\theta}}})$ for some evaluation function $f$. For instance, in our experiments, $f$ is typically test loss or a test prediction. To show that the RIF and NS estimates are close, we require our evaluation function $f$ to have bounded gradients:
@@ 4 Related Work
The seminal work KL (17) introduced the modern use of influence functions to study the relationship between training data and model behavior in modern machine learning. Ensuing works BNL+ (22); BPF (21); GBA+ (23); FZ (20) study influence functions for neural networks, and use them as a tool to study and interpret model behavior. GJB (19); BYF (20) propose second and higher-order approximations to leave-one-out and leave-$T$-out effects, but these approximations sacrifice linearity and efficiency. Many applications of influence functions have appeared recently, e.g., machine unlearning GGHVDM (19); SAKS (21); SW (22), data valuation JDW+ (19), robustness quantification SS (19), and fairness LL (22). To scale influence functions up to very large models and datasets, where Hessian inversion becomes infeasible, several works develop sketching/random projection techniques to approximate influence functions, e.g., WCZ+ (16); PGI+ (23); SZVS (22).
Newton-step approximations to the leave-1-out error have been studied since at least 1981 Pre (81). Cross-validation is an especially important application RM (18); WKM (20). Even though Newton-step approximations are long-studied, to the best of our knowledge, the idea behind RIF – adding leave-1-out Newton step approximations to get a leave-$T$-out approximation – is novel.
@@ Acknowledgments
This work was supported by NSF Award No. 2238080.
Compute Resources
All experiments were conducted on a server equipped with 64GB RAM, 2 IBM POWER9 CPU cores, and 4 NVIDIA Tesla V100 SXM2 GPUs (each with 32GB memory). Compute resources were not a bottleneck for our work. The total wall-clock time for all experiments reported in the paper was under 100 hours.
@@
Newton $\textup{Newton Step}=\left(\nabla f\right)^{\top}\left({\mathbf{H}}-\sum_{j=1}^{n}w_{j}{\mathbf{H}}_{j}\right)^{-1}\sum_{i=1}^{n}w_{i}{\mathbf{g}}_{i},$
@@
    The higher order terms$\left(\nabla f\right)^{\top}{\mathbf{H}}^{-1}\left({\mathbf{H}}-{\mathbf{H}}_{\mathbf{w}}\right){\mathbf{H}}_{\mathbf{w}}^{-1}\left({\mathbf{H}}_{{\mathbf{w}}^{(i)}}-{\mathbf{H}}_{\mathbf{w}}\right){\mathbf{H}}_{{\mathbf{w}}^{(i)}}^{-1}{\mathbf{g}}_{i}$.
@@
|Newton ${\left\lvert{\textup{Newton Step}-\textup{RIF}}\right\rvert}={\left\lvert{\sum_{i=1}^{n}w_{i}\left(\nabla f\right)^{\top}\left({\mathbf{H}}_{\mathbf{w}}^{-1}-{\mathbf{H}}_{{\mathbf{w}}^{(i)}}^{-1}\right){\mathbf{g}}_{i}}\right\rvert}\leq k^{2}\eta\left(1+2C_{R}\right)\left(\varepsilon+C_{R}C_{\ell}\delta\right)$
@@
$C({\mathbf{M}})C({\mathbf{N}})=C({\mathbf{L}}),\quad\text{where }{\mathbf{L}}_{ij}=\sum_{q,r}{\mathbf{M}}_{iq}\cdot\sqrt{{\mathbf{A}}_{q}}{\mathbf{H}}^{-1}\sqrt{{\mathbf{A}}_{r}}\cdot{\mathbf{N}}_{rj}.$
@@ B.1 Analysis of [22]
The matrix ${\mathbf{D}}(w)$ has eigenvalues between 00 and $\sigma_{\max}/\lambda$. The residual term $\mathrm{Err}_{f,3}(w)$ captures the error due to third-order derivatives and is bounded by
@@ C.1 Model Training
For the DogFish and Enron datasets also considered by Koh et al., we used the same $L_{2}$ regularization parameter, and for all new datasets, we set the regularization to $51E-$ 5.
@@ C.3 Datasets and Embeddings
We embed these audio samples using last-layer embeddings of the OpenL3 python library [6]. This produces $d=512$ dimensional embeddings, and we separate them into train and test samples using a random $80-20$ train-test split.

### v2

@@ Abstract
How does the training data affect a model’s behavior? This is the question we seek to answer with data attribution. The leading practical approaches to data attribution are based on influence functions (IF). IFs utilize a first-order Taylor approximation to efficiently predict the effect of removing a set of samples from the training set without retraining the model, and are used in a wide variety of machine learning applications. However, especially in the high-dimensional regime (# params $\geq\Omega($# samples))), they are often imprecise and tend to underestimate the effect of sample removals, even for simple models such as logistic regression. We present rescaled influence functions (RIF) – a tool for data attribution which can be used as a drop-in replacement for influence functions, with little computational overhead but significant improvement in accuracy. We compare IF and RIF on a range of real-world datasets, showing that RIFs offer significantly better predictions in practice, and present a theoretical analysis explaining this improvement. Finally, we present a simple class of data poisoning attacks that would fool IF-based detections but would be detected by RIF.
@@ 1 Introduction
The ability to quickly and accurately predict a leave-$T$-out (LTO) effect, or to search for subsets producing a large leave-out effect, unlocks extensive capabilities from classical statistical inference to modern machine learning. For example, the jackknife, leave-$T$-out cross-validation, and bootstrap are all widely used to quantify uncertainty and estimate generalization error or confidence intervals, and all rely on the ability to quickly estimate LTO effects Efr (92); GSL+ (19); Jae (72). Machine learning has seen an explosion of applications of data attribution, for dataset curation KL (17); KATL (19), explainability KATL (19); GBA+ (23), crafting and detection of data poisoning attacks EIC+ (25); KSL (22); SS (19), machine unlearning SAKS (21); GGHVDM (19); IASCZ (21), credit attribution JDW+ (19); GZ (19), bias detection BAHAZ (19), and more.
Ascertaining the ground truth leave-$T$-out effect in general requires a full retrain of a model for each $T$ of interest, which is computationally intractable in all but the simplest settings. Consequently, approximations to the leave-$T$-out effect are widely used. Key desiderata for such approximations are (1) accuracy, (2) computational efficiency even for large-scale models, and (3) additivity: the predicted effect of removing $T$ should be the sum of predicted effects of removing each element of $T$ individually. Additivity enables another important capability: search for the subset $T$ of a given size with the greatest predicted effect according to a given metric, by taking the $k$ training data points with largest predicted leave-one-out (LOO) effects BGM (20); IPE+ (22); HBN+ (24).
@@
We study a simple and fast-to-compute modification of the influence function, which we term the rescaled influence function (RIF). RIFs improve accuracy by incorporating a limited amount of higher-order information about the change in model parameters from sample removal, but retain the additivity and in many settings also the computational efficiency of IFs. We show via experiments and theoretical analysis that RIFs are accurate for data attribution in overparameterized models where IFs struggle. Like IFs, RIFs are model and task agnostic, meaning that they can be applied to any empirical risk minimization-based training method with smooth losses, and they can estimate the leave-$T$-out effect according to any (smooth) measure of change to model parameters. We therefore advocate using RIFs as a drop-in replacement for IFs across data attribution applications.
@@ 1.1 Influence Functions, Newton Steps, and Rescaled Influence Functions
We now introduce the rescaled influence function formally. Suppose that $\{(x_{i},y_{i})\}_{i\in[n]}$ is a training data set, $\Theta\subseteq\mathbb{R}^{d}$ is a class of models, and $\ell(x,y,\theta)$ is a twice-differentiable loss function; $\ell$ may include a regularizer. For simplicity, we imagine that $\ell$ is convex, although the definition of RIFs can be extended to the non-convex case. Let $\hat{\theta}=\arg\min_{\theta\in\Theta}\sum_{i\leq n}\ell(x_{i},y_{i},\theta)$ be the empirical loss minimizer (or, in the non-convex setting, any local minimum of the empirical loss).
@@
[…]
@@
[…]
@@
[…]
@@
The computational overhead of RIF compared to IF depends in general on the cost of computing the $n$ leave-one-out Hessian inversions – once these are obtained, no fresh Hessian inversion is needed to compute $\hat{\theta}_{\textup{RIF},T}$ for any $T$. RIF is especially attractive in generalized linear models and neural networks with a ReLU activation function, where $\textup{RIF}_{i}$ can be obtained from $\text{IF}_{i}$ by multiplying by a rescale factor $(1-h_{i})^{-1}$, where $h_{i}$ is a (generalized) leverage score associated to the $i$-th sample, which can be computed via a single matrix-vector product with $H^{-1}$. Thus, for generalized linear models, no additional Hessian inversion is needed. For example, in logistic regression, the formula for $\textup{RIF}_{i}$ uses the rescaling $(1-h_{i})^{-1}$, where $h_{i}=\hat{y_{i}}(1-\hat{y_{i}})\cdot x_{i}^{\top}H^{-1}x_{i}$; here $\hat{y_{i}}\in[0,1]$ is the logistic predicted label of the $i$-th sample according to $\hat{\theta}$.
Beyond generalized linear models and ReLU neural networks, whenever each sample makes a low-rank contribution the Hessian, the $n$ leave-one-out Hessian inversions can be computed quickly via the Sherman-Morrison/Woodbury formula. In all of our experiments, the running time overhead to compute RIF is negligible (see Table 2).
@@
We note that the idea of summing over estimates of leave-one-out effects to estimate the leave-$T$-out effect is not new, and has been a central component of many previous data models IPE+ (22). In their seminal TRAK paper, Park et al. separately consider both the idea of combining LOO effects additively PGI+23a [Definition 2.3] and the idea of using a Newton step to estimate LOO effects of a logistic regression PGI+23a [Definition 3.1] but do not explicitly combine the rescaling effect in their estimator except to note that the rescaling correction has little to no effect in their setting.
A similar approach that has been the focus of recent research is the Additive-One-Exact data model, which estimates the LTO effect by summing over the exact LOO effects. This data model was introduced by Kuschnig et al. KZCC (21) and further analyzed by Huang et al. HBN+ (24). Both Kuschnig et al. and Huang et al. study the accuracy of this method for identifying sets of highly influential samples in ordinary least squares (OLS) regressions. Moreover, Huang et al. also note that because a single Newton step is equivalent to a full retrain for the case of OLS, a natural extension of the Additive-One-Exact data model is to sum over the single-Newton step attributions of the individual samples HBN+ (24)[Appendix C.2]. But, to the best of our knowledge, no prior work offers quantitative experimental or theoretical comparisons between RIF and other data attribution methods in the high-dimensional settings where the differences we study emerge, or beyond the case of OLS where it is equivalent to the Additive-One-Exact model. [^2] We are grateful to Tamara Broderick and Jenny Huang for making us aware of these prior works via personal communication.
@@ 3 Theoretical Results
[…]
@@
[…]
@@ 4 Related Work
The seminal work KL (17) introduced the modern use of influence functions to study the relationship between training data and model behavior in modern machine learning. Ensuing works BNL+ (22); BPF (21); GBA+ (23); FZ (20) study influence functions for neural networks, and use them as a tool to study and interpret model behavior. GJB (19); BYF (20) propose second and higher-order approximations to leave-one-out and leave-$T$-out effects, but these approximations sacrifice linearity and efficiency. Many applications of influence functions have appeared recently, e.g., machine unlearning GGHVDM (19); SAKS (21); SW (22), data valuation JDW+ (19), robustness quantification SS (19), and fairness LL (22). To scale influence functions up to very large models and datasets, where Hessian inversion becomes infeasible, several works develop sketching/random projection techniques to approximate influence functions, e.g., WCZ+ (16); PGI+23b ; SZVS (22).
@@
Newton-step approximations to the leave-1-out error have been studied since at least 1981 Pre (81). Cross-validation is an especially important application RM (18); WKM (20). Additionally, several recent works consider data models that additively combine estimates of leave-one-out effects to compute a leave-$T$-out effect KZCC (21); IPE+ (22); PGI+23a ; HBN+ (24). However, to the best of our knowledge no previous work provides an empirical or theoretical evaluation of the RIF method beyond low-dimensional least-squares regression.
@@ Conclusion
Compute Resources
All experiments were conducted on a server equipped with 64GB RAM, 2 IBM POWER9 CPU cores, and 4 NVIDIA Tesla V100 SXM2 GPUs (each with 32GB memory).
Table 2 details the computational cost of training the base models and computing their IF and RIF data attribution. Another major computational overhead was in retraining the model to obtain ground-truth values for the retrain effect. Despite this, compute resources were not a bottleneck for our work. The total wall-clock time for all experiments reported in the paper was under 100 hours.
Dataset Training Hessian Inversion Influence Rescaling
ESC50 1.8 s 0.056 s 0.0005 s 0.051 s 0.0033 s (0.2%)
CatDog 76 s 4.9 s 0.010 s 4.8 s 0.087 s (0.1%)
AutoTruck 48 s 4.9 s 0.0094 s 4.8 s 0.087 s (0.2%)
DogFish 0.43 s 0.92 s 0.0095 s 0.89 s 0.015 s (0.7%)
Enron 6.7 s 15 s 0.065 s 15 s 0.095 s (0.3%)
IMDB (n=16d) 20 s 0.92 s 0.0012 s 0.87 s 0.044 s (0.2%)
Table 2: Comparison of runtime components across datasets. The Rescaling step consistently added negligible overhead across all experiments.
@@ Acknowledgments
We would like to thank Jenny Y. Huang, David R. Burt, Yunyi Shen, Tin D. Nguyen, Vishwak Srinivasan, and Tamara Broderick for helpful conversations. This work was supported by NSF Award No. 2238080 and CSAIL Alliances.
@@
$\textup{Newton Step}=\left(\nabla f\right)^{\top}\left({\mathbf{H}}-\sum_{j=1}^{n}w_{j}{\mathbf{H}}_{j}\right)^{-1}\sum_{i=1}^{n}w_{i}{\mathbf{g}}_{i},$
@@
[…]
@@
${\left\lvert{\textup{Newton Step}-\textup{RIF}}\right\rvert}={\left\lvert{\sum_{i=1}^{n}w_{i}\left(\nabla f\right)^{\top}\left({\mathbf{H}}_{\mathbf{w}}^{-1}-{\mathbf{H}}_{{\mathbf{w}}^{(i)}}^{-1}\right){\mathbf{g}}_{i}}\right\rvert}\leq k^{2}\eta\left(1+2C_{R}\right)\left(\varepsilon+C_{R}C_{\ell}\delta\right)$
@@
[…]
@@ B.1 Analysis of [23]
The matrix ${\mathbf{D}}(w)$ has eigenvalues between 0 and $\sigma_{\max}/\lambda$. The residual term $\mathrm{Err}_{f,3}(w)$ captures the error due to third-order derivatives and is bounded by
@@ C.1 Model Training
For the DogFish and Enron datasets also considered by Koh et al., we used the same $L_{2}$ regularization parameter, and for all new datasets, we set the regularization to 1E51E-5.
@@ C.3 Datasets and Embeddings
We embed these audio samples using last-layer embeddings of the OpenL3 python library [6]. This produces $d=512$ dimensional embeddings, and we separate them into train and test samples using a random 802080-20 train-test split.